A body of mass 2kg is placed on a rough inclined plane whose angle of inclination is 30 degree. Find the resultant force acting down on the slope and acceleration if a frictional force of 5N acts between the body and the inclined surface.
1
Expert's answer
2019-07-05T14:15:38-0400

The resultant force acting down on the body along a slope


"F=mg \sin\theta=2\:\rm{kg}\times 9.8\:\rm{N/kg}\times\sin 30^{\circ}=9.8\:\rm{N}"

The acceleration of the body

"a=\frac{F-F_{\rm{fric}}}{m}=\frac{9.8\:\rm{N}-5\:\rm{N}}{2\:\rm{kg}}=2.4\:\rm{m/s^2}"